Use the given standard enthalpies of formation to determine the heat of reaction of the following reaction:

2 C5H10O(g) + 19 O2(g) => 10 CO2(g) + 10 H2O(g)

        ΔHof C5H10O(g) = -232.11 kJ/mol

Δ Hof CO2(g) = -393.5 kJ/mol

        ΔHof H2O(g) = -241.8 kJ/mol
Answer #1

8CshoO+ 110, 6) 6 , + TO HO (0) ΔΗ - ΣΔ» πιει ΣΑ noseis- = (ox-3135+los-241 - [ax-232) – –3535 -24+ +6+2) KT ΔΑ = -5888 - 16z